# Learning Agility Adaptation for Flight in Clutter

### v1

@@ Abstract
Animals learn to adapt agility of their movements to their capabilities and the environment they operate in. Mobile robots should also demonstrate this ability to combine agility and safety. The aim of this work is to endow flight vehicles with the ability of agility adaptation in prior unknown and partially observable cluttered environments. We propose a hierarchical learning and planning framework where we utilize both trial and error to comprehensively learn an agility policy with the vehicle’s observation as the input, and well-established methods of model-based trajectory generation. Technically, we use online model-free reinforcement learning and a pre-training-fine-tuning reward scheme to obtain the deployable policy. The statistical results in simulation demonstrate the advantages of our method over the constant agility baselines and an alternative method in terms of flight efficiency and safety. In particular, the policy leads to intelligent behaviors, such as perception awareness, which distinguish it from other approaches. By deploying the policy to hardware, we verify that these advantages can be brought to the real world.
@@ I Introduction
Animals seldom move at their maximum speeds due to limited sensory, reaction, or motor capabilities. For instance, some animals, such as birds, enhance the resolution of spatial perception by slowing down while foraging [1]. Cheetahs almost never chase their prey at full speed due to the difficulties of sharp turns or footing maintenance [2]. Such compromising behaviors are especially likely to occur in constrained environments [3, 4], where animals regulate their agility to ensure safety considering their limited capabilities, e.g., budgerigars fly at a low speed to ensure collision-free crossing of narrow gaps [5].
@@
The same goes for mobile robots. The limited sensory update frequency, time-consuming decision-making, and imperfect motor control capabilities, to name a few, inherently limit the allowed agility of them that satisfies safety regards. This agility-safety trade-off is always present, no matter how far a hardware or algorithmic system has evolved. Therefore, like animals, robots should be able to adaptively regulate their level of agility based on an integrated cognition relating self-awareness, e.g., of their own capability limitations, and other-awareness, e.g., of the external environments [6].
@@
This paper focuses on flight vehicles, where some agile behaviors have been preliminarily achieved [7, 8, 9, 10]. These advances are made possible by the considerable development of model-based trajectory planners [11, 7, 12] and low-level controller [8, 13, 14] in the last decades, and the recent application of model-free learning techniques [9, 10]. However, most of the existing planning and control schemes leave the task of determining the agility level for the user, which is conservatively set to constant at deployment time [7, 15]. While some works utilize reinforcement learning (RL) to learn a policy that directly outputs low-level commands with implicit agility adaptations [9, 10], the success of such approaches occurs for the time being only in prior known and global observable environments. In contrast, this paper considers the problem of safe flight in unknown, partially observed, and cluttered environments.
@@
Since expert policy for agility adaptation is unavailable, an intuitive idea is to use online RL to learn a policy from scratch to enable naturally embedded agility adaptation, as implemented by a concurrent work [16]. However, such an approach is expensive to learn a policy, and at this stage can only be deployed for simpler scenarios than a state-of-the-art model-based trajectory planner can handle [7, 17]. Instead, in this paper, we take advantage of the insights gained over the past decades in classical trajectory generation and tracking frameworks, which are favored due to their formal safety guarantee and generalizability that is not available by policies learned from scratch.
@@
In particular, we decompose the policy into a hierarchical one, where the outer-loop policy decides the current level of agility and can be effectively learned with model-free RL. The inner-loop policy, which is conditioned on the output of the outer-loop policy, generates the trajectory for execution. Fortunately, modern model-based trajectory planners [7, 17, 15], despite their requirement for a pre-determined agility level, have evolved considerably to serve as a near-optimum to the inner-loop policy. Similar hierarchical frameworks that combine RL and model-based optimization can be found in problems such as model predictive control (MPC) [18, 19], legged locomotion [20, 21], and automated machine learning (AutoML) [22].
Contribution. This paper contributes a hierarchical learning and planning approach for adaptive agile and safe flight in unknown cluttered environments. The policy that generates trajectories for execution is decomposed, where an agility sub-policy optimizes the whole policy by learning.
Technically, we employ a pre-training-fine-tuning rewarding and training scheme to overcome the sparsity failure modes challenge. The first stage of reward imposes human knowledge to facilitate training and pave the way for the second stage of training. The second stage of reward, on the other hand, directly reflects our objective goals to combine agility and safety.
We validate the significance of agility adaptation in simulation, where the proposed approach outperforms both constant agility baselines and an alternative approach [23]. The policy can be directly deployed in the real world, leading to adaptive and safe flight in clutter, as shown in Fig. LABEL:fig:head.
@@ II-A Adaptive Motion Planning
The works that have the most similar motivation to us are [23, 24, 25], which endeavor to alter agility for flight with model-based motion planners. In [23, 25], the authors impose the desired adaptive behavior by incorporating the velocity into trajectory planning via handcrafted cost functions. However, these cost functions are not expressive enough to comprehensively address the problem and inevitably lead to inflexible behavior. Methodologically, while we believe that powerful modeling-based approaches can be raised through more generalized and accurate formulations [26] and specialized computational techniques, learning-based approaches, if feasible, can serve as a simpler alternative. Zhou et al. [24] employ an online learning method based on Beyasian optimization to decide hyperparameter for trajectory planners. However, the painful convergence speed of Beyasian optimization makes their method incapable of adapting to a rapidly updating observation. We believe that an offline training mode, where the policy is determined before deployment, is more appropriate for the goal. Richter et al. [27, 28] formalize the problem of planning in an unknown environment and indicate the fundamental intractability in this formulation is on the belief that must capture the distribution of environments. Instead of directly operating on this distribution, the authors propose to learn a collision probability based on hand-coded features, while by striking out artificial features, our method is expected to achieve more flexible pattern recognition by jointly training perception and action modules.
As mentioned in section I, a work [16] released within weeks of our submission coincidentally studies agility-adaptive flight, but follows a different idea by training a policy directly outputting acceleration commands. In this approach, agility level is not explicitly incorporated into the action space. A specialized latent space used to maintain historical information is the key for the policy, which is achieved by a model-based mapping algorithm [29] in our framework. Our approach is tested in clutters (see Fig. LABEL:fig:head, Fig.5, and Fig. 10) that are much more complex compared to those in [16]. The unique success of our approach stems not only from the learned policy, but also the generalizability and safety of the model-based trajectory planner [17].
@@ II-B Combining Learning and Model-based Planner for Navigation in Cluttered Environments
Recently, some works enhance traditional planning and control frameworks with learned policy to achieve more robust or efficient performance. Previous works [30, 15] employ imitation learning to direct predict local waypoints from visual input without geometric mapping for perceptive navigation. RL is also used to find an optimal policy [31] or a value estimator [32, 33] as additional modules for navigation without manual labeling. The most similar of these works to ours in terms of technical pipelines are [34, 35], which employ RL to regularize the parameters for a classical local planning algorithm, Dynamic Window Approach (DWA). However, our problem has distinct technical differences from these works. First, since we are concerned with combining agility and safety, we have to deal with sparse collision rewards. In many works, collision penalties can be densified by relaxing them as distance to the nearby obstacles. However, on the one hand, in this work, the spatial distribution of trajectories generated by the planner is almost independent of the output of RL policy, so penalizing distance-related metrics is not causally meaningful. On the other hand, especially in regions with densely distributed obstacles, penalizing distance is not a good reflection of our expectation on the behavior of vehicle and may introduce a short-sightedness. Second, the vehicle has to deal with limited field of view (FOV), rather than the near omnidirectional perception in [34, 35], which requires it to behave in a perception-aware manner. Moreover, while the above works are about ground robots navigating with low speed, our work contributes to agile flight systems, which requires the RL policy to cooperate with 3D perception and a much more complex modern planner backbone. Therefore, our work has a unique application contribution.
@@ III Problem Formulation
  • States $S$. The states space is divided into two parts, one is the controllable states $S^{c}$ which consists of the states of the vehicle such as $\boldsymbol{p}$ (position), $\boldsymbol{v}$ (velocity), $\boldsymbol{a}$ (acceleration), etc. The other part of $S$ is the environment $S^{e}$. In the context of our problem, $S^{e}$ can be described as the occupancy of the space, i.e., $s^{e}=\{0,1\}^{n}$.
  • Actions $A$. Actions can be thought of control commands in a general sense such as the output of trajectory planning. For the class of flight vehicles considered in this paper, the specific form of the trajectory is reviewed in section IV-A.
  • Conditional transition probabilities $T$. We assume that the dynamics of vehicle is known and deterministic, while the environment is static, i.e., $p\left(s^{e}|s^{e},a_{t}\right)=1$.
  • The reward function $R$. As the topic of this paper suggests, our reward function is defined as maximizing agility and minimizing collision loss, where agility is defined as the velocity component of the trajectory
  • Observations $\varOmega$. Observations include the state estimates of the vehicle, which is assumed to be perfect, and the partial observation of the environment $s^{e}$. The latter is limited by the FOV of the sensor (in this paper, the depth camera) and occlusion by obstacles.
@@
The action at each moment is generated by a policy conditioned on the current observation and goal of navigation $g\in G$, i.e., $a_{t}\sim\pi\left(\cdot|o_{t},g\right)$. Our goal is to find the policy that maximize the expected cumulative return over time, i.e.,
@@ IV Framework
Our framework involves a learned policy, which is donated as the agility policy $\pi^{\dagger}$, a model-based trajectory planner, and the standard perception and low-level control ends. The agility policy and model-based trajectory planner form a hierarchical policy to optimize (1).
@@ IV-A Model-based Trajectory Planner
$\min_{u\left(t\right)}\int_{0}^{T}{J\left(u\left(t\right)\right)+\rho\left(T\right)},\vspace{-0.1cm}$ (2a)
@@
$\mathcal{H}\left(u\left(t\right)\right)=\boldsymbol{0},\ \ \forall t\in\left[0,T\right],$ (2d)
where $u\left(t\right)$ is the (high-level) planning quantities, $\rho:\left[\left.0,\infty\right)\right.\mapsto\left[0,\infty\right]$ the time regularization term, $\mathcal{F}$ the collision free region, $\mathcal{G}\left(\cdot\right)$ the dynamic constraints imposed by the physical limitation of the vehicle or users, and $\mathcal{H}\left(\cdot\right)$ summarizes the equality relations between the planning quantities and the states or control commands [36]. A widely used $u$ for underactuated multicopters consists of the flat-output space $z=\left(p_{x},p_{y},p_{z},\psi\right)^{\textnormal{T}}$ and the finite derivatives of it [11], where $(p_{x},p_{y},p_{z})^{\textnormal{T}}$ is the translation of the center of gravity and $\psi$ the yaw angle of the vehicle.
In the problem of navigation in cluttered environments, an inequality in (2c) represents the level of agility, such as $\left\|\boldsymbol{v}\right\|\leq\bar{v}$ or $\boldsymbol{v}\preceq\bar{\boldsymbol{v}}$, where $\bar{v}$ and $\bar{\boldsymbol{v}}$ are the imposed restriction on agility [7, 17]. For the generality of our description, we write such constraints uniformly as
$h\left(\boldsymbol{v}\right)\preceq v^{\dagger}\vspace{-0.1cm}$ (3)
@@ IV-B Hierarchical Policy Optimization
Traditionally, $v^{\dagger}$ is conservatively determined in advance based on the user’s observations of the environment to ensure safety. However, such an approach does not reasonably take advantage of the perception capabilities of the sensors, the performance of the algorithmic system, and the maneuverability of the vehicle. Instead, jointly optimizing the agility and trajectory to optimize (1) can further unlock the potential of the system.
We define a decomposition of the policy augmented with an auxiliary action $v^{\dagger}$
@@
where the policy $\pi^{\prime}$ is decomposed as two hierarchical policies $\pi^{\prime\prime}$ and $\pi^{\dagger}$, the former of which is conditioned on the auxiliary action while the latter decides it. The action space of the original POMDP is augmented as a result of the introduction of an auxiliary action, i.e., $A^{\prime}=A\times V^{\dagger}$. However, no other part of the POMDP is changed, so we can define a hierarchical policy optimization problem
$\begin{split}\max_{\pi^{\prime\prime},\pi^{\dagger}}\mathbb{E}_{a_{t}\sim\pi^{\prime\prime}(\cdot|v^{\dagger}_{t},o_{t},g),v^{\dagger}_{t}\sim\pi^{\dagger}\left(\cdot|o_{t},g\right)}\left[\sum\nolimits_{t}{\gamma^{t}r\left(s_{t},a_{t},g\right)}\right]=\\ \max_{\pi^{\dagger}}\max_{\pi^{\prime\prime}}\mathbb{E}_{a_{t}\sim\pi^{\prime\prime}(\cdot|v^{\dagger}_{t},o_{t},g),v^{\dagger}_{t}\sim\pi^{\dagger}(\cdot|o_{t},g)}\left[\sum\nolimits_{t}{\gamma^{t}r(s_{t},a_{t},g)}\right],\end{split}$ (5)
where, for more concise writing, we omit the state and goal distributions in (1). This problem is equivalent to (1) in the sense that if $\pi$ and $\pi^{\prime}$ impose the same distribution on $a$ when conditioned on any $o\in O$ and $g\in G$, the objective function values of (1) and (5) are consistent.

### v2

# Learning Speed Adaptation for Flight in Clutter

@@ Abstract
Animals learn to adapt speed of their movements to their capabilities and the environment they observe. Mobile robots should also demonstrate this ability to trade-off aggressiveness and safety for efficiently accomplishing tasks. The aim of this work is to endow flight vehicles with the ability of speed adaptation in prior unknown and partially observable cluttered environments. We propose a hierarchical learning and planning framework where we utilize both well-established methods of model-based trajectory generation and trial-and-error that comprehensively learns a policy to dynamically configure the speed constraint. Technically, we use online reinforcement learning to obtain the deployable policy. The statistical results in simulation demonstrate the advantages of our method over the constant speed constraint baselines and an alternative method in terms of flight efficiency and safety. In particular, the policy behaves perception awareness, which distinguish it from alternative approaches. By deploying the policy to hardware, we verify that these advantages can be brought to the real world.
@@ I Introduction
Animals seldom move at their maximum speeds due to limited sensory, reaction, or motor capabilities. For instance, some animals, such as birds, enhance the resolution of spatial perception by slowing down while foraging [1]. Cheetahs almost never chase their prey at full speed due to the difficulties of sharp turns or footing maintenance [2]. Such compromising behaviors are especially likely to occur in constrained environments [3], where animals regulate their speed to ensure safety considering their limited capabilities, e.g., budgerigars fly at a low speed to ensure collision-free crossing of narrow gaps [4].
@@
The same goes for mobile robots. The limited sensory update frequency, time-consuming decision-making, and imperfect motor control capabilities, to name a few, inherently limit the allowed speed of their movements that satisfies safety regards. This efficiency-safety trade-off is always present, no matter how far a hardware or algorithmic system has evolved. Therefore, like animals, robots should be able to adaptively regulate their speed of movements based on an integrated cognition relating self-awareness, e.g., of their own capability limitations, and other-awareness, e.g., of the external environments [5].
@@
This paper focuses on flight vehicles, where some agile behaviors have been preliminarily achieved [6, 7, 8, 9]. These advances are made possible by the considerable development of model-based trajectory planners [6, 10] and controllers [7, 11] in the last decades, and the recent application of model-free learning techniques [8, 9]. However, most of the existing planning and control schemes leave the task of determining the speed constraint for the user, which is conservatively set to constant at deployment time [6, 12]. While some works utilize reinforcement learning (RL) to learn a policy that directly outputs low-level commands with implicit speed adaptations [8, 9], the success of these works occurs for the time being only in prior known environments. In contrast, this paper considers the problem of safe flight in unknown, partially observed, and cluttered environments.
@@
An intuitive idea is to trial and error, learning a policy from scratch to enable naturally embedded speed adaptation according to the inherent limitations of the system, as implemented by a concurrent work [13]. However, such an approach is expensive to learn a deployable policy, and at this stage can only be deployed for simpler scenarios than a state-of-the-art model-based trajectory planner can handle [6, 12]. Instead, in this paper, we take advantage of the insights gained over the past decades in classical trajectory generation and tracking frameworks, which are favored due to their formal safety guarantee and generalizability that is not available by policies learned from scratch.
@@
In particular, we decompose the policy into a hierarchical one, where the outer-loop policy dynamically configures the speed constraint and can be effectively learned with online RL. The inner-loop policy, which is conditioned on the output of the outer-loop policy, generates the trajectory for execution. Fortunately, modern model-based trajectory planners [6, 12, 14], despite their requirement for a pre-determined constant speed constraint, have evolved considerably to serve as a near-optimum to the inner-loop policy in the scenarios they are specifically designed for. Similar hierarchical frameworks that combine experience-based learning and model-based optimization can be found in problems such as model predictive control (MPC) [15, 16, 17] and legged locomotion [18, 19].
Our main contribution is a system that hierarchically combines the model-based trajectory planner and a learned outer-loop policy to enable aggressiveness-adaptive flight in clutter. This approach is benchmarked to outperform baselines with constant constraints and an alternative approach [20]. We also demonstrate the system in real-world scenarios, including a complex natural clutter shown in Fig. LABEL:fig:head, exhibiting aggressive but safe flight in the wild. One of the crucial technical designs to effectively learn the outer-loop policy is a two-stage reward scheme that is employed to overcome the challenge of stochasticity and sparsity posed by the early-termination penalty.
@@ II-A Adaptive Motion Planning
The works that have the most similar motivation to us are [20, 21, 22], which endeavor to alter flying speed with model-based motion planners. In [20, 22], the authors impose the desired adaptive behavior by incorporating the velocity into trajectory planning via handcrafted cost functions. However, these cost functions are not expressive enough to comprehensively address the problem and inevitably lead to inflexible behavior. Methodologically, while we believe that powerful modeling-based approaches can be raised through more generalized and accurate formulations and specialized computational techniques, learning-based approaches, if feasible, can serve as a simpler alternative. Zhou et al. [21] employ an online learning method based on Beyasian optimization to decide hyperparameter for trajectory planners. However, the painful convergence speed of Beyasian optimization makes their method incapable of adapting to a rapidly updating observation. We believe that an offline training mode, where the policy is determined before deployment, is more appropriate for the goal. Richter et al. [23] formalize the problem of planning in an unknown environment and indicate the fundamental intractability in this formulation is on the belief that must capture the distribution of environments. Instead of directly operating on this distribution, the authors propose to learn a collision probability based on hand-coded features, while by striking out artificial features, our method is expected to achieve more flexible pattern recognition by jointly training perception and action modules.
As mentioned in section I, a work [13] released within weeks of our submission coincidentally studies speed-adaptive flight, but follows a different idea by training a policy directly outputting acceleration commands. In this approach, speed constraint is not explicitly incorporated into the action space. A specialized latent space used to maintain historical information is the key for the policy, which is achieved by a model-based mapping algorithm [24] in our framework. Our approach is tested in clutter (see Fig. LABEL:fig:head, Fig. 4, and Fig. 9) that are much more complex compared to those in [13]. The unique success of our approach stems not only from the learned policy, but also the generalizability and safety of the model-based trajectory planner [12].
@@ II-B Combining Learning and Model-based Planner for Navigation in Cluttered Environments
Recently, some works enhance traditional planning and control frameworks with learned modules to achieve more robust or efficient performance. Previous works [25, 14] employ imitation learning to directly predict local waypoints from visual input without geometric mapping for perceptive navigation. RL is also used to find an optimal policy [26] or a value estimator [27] as additional modules for navigation without manual labeling. The most similar of these works to ours in terms of technical pipelines are [28, 29], which employ RL to regularize hyperparameters for a classical local planning algorithm, Dynamic Window Approach (DWA). In these works, collision penalties can be densified by relaxing them as distance to the nearby obstacles. However, in this work, the spatial distribution of trajectories generated by the planner is almost independent of the output of the learned policy, so penalizing distance-related metrics is not causally meaningful, for which we employ a two-stage reward scheme. Moreover, the vehicle in our setup also has to deal with the limited field of view (FOV), rather than the near omnidirectional perception in [28, 29], which requires it to behave in a perception-aware manner. We explicit enable this by designing a local map representation (see section V-B). While the above works are about ground robots navigating with low speed, our work contributes to agile flight systems, which requires the RL policy to cooperate with 3D perception and a much more complex modern planner backbone. Therefore, our work has a unique application contribution.
@@ III Problem Formulation
  • States $S$. The states space is divided into two parts, one is the controllable states $S^{c}$ which consists of the states of the vehicle. The other part of $S$ is the environment $S^{e}$. In the context of our problem, $S^{e}$ can be described as the occupancy of the space, i.e., $s^{e}=\{0,1\}^{n}$.
  • Actions $A$. Actions can be thought of as control commands in a general sense, which can be outputs of the trajectory planner or the actuator of the vehicle.
  • Conditional transition probabilities $T$. According to the definition of $S$, the transition probability is also divided into two independent components, where we impose no assumption to the state transition of $s^{c}$, while assuming $p(s^{e}|s^{e},a_{t})=1$, i.e., the environment $S^{e}$ is static.
  • The reward function $R$. As the topic of this paper suggests, our reward function is defined as maximizing movement speed and minimizing collision loss.
  • Observations $\varOmega$. Observations include the state estimates of the vehicle, which is assumed to be perfect, and the partial observation of the environment $s^{e}$. The latter is limited by the FOV of the external sensor (in this paper, the depth camera) and occlusion by obstacles.
@@
The action at each moment can be viewed as generated by a policy conditioned on the current observation and goal of navigation $g\in G$, i.e., $a_{t}\sim\pi\left(\cdot|o_{t},g\right)$. The goal is to find the policy that maximizes the expected cumulative return,
@@ IV Framework
Our framework involves a learned policy, which is donated as $\pi^{\dagger}$, a model-based trajectory planner, and the standard perception and low-level control ends. The policy $\pi^{\dagger}$ and model-based trajectory planner form a hierarchical policy to optimize (1).
@@ IV-A Model-based Trajectory Planner
$\min_{u\left(t\right),T}\int_{0}^{T}{J\left(u\left(t\right)\right)+\rho\left(T\right)},$ (2a)
@@
$\mathcal{H}\left(u\left(t\right)\right)=\boldsymbol{0},\ \ \forall t\in\left[0,T\right],\vspace{-0.1cm}$ (2d)
where $u\left(t\right)$ is the planning quantities, $\rho:\left[\left.0,\infty\right)\right.\mapsto\left[0,\infty\right]$ the time regularization term, $\mathcal{F}$ the collision free region, $\mathcal{G}\left(\cdot\right)$ the dynamic constraints imposed by the physical limitation of the vehicle or users, and $\mathcal{H}\left(\cdot\right)$ summarizes the equality relations between the planning quantities and the states or control commands [30].
In the problem of navigation in cluttered environments, an inequality in (2c) represents the level of aggressiveness on speed of vehicle $\boldsymbol{v}$, such as $\left\|\boldsymbol{v}\right\|\leq\bar{v}$ or $\boldsymbol{v}\preceq\bar{\boldsymbol{v}}$ [6, 12]. For the generality of our description, we write such constraints uniformly as
$h\left(\boldsymbol{v}\right)\preceq v^{\dagger}$ (3)
@@ IV-B Hierarchical Policy Optimization
Traditionally, $v^{\dagger}$ is conservatively determined in advance based on the user’s observations of the environment to ensure safety. However, such an approach does not reasonably take advantage of the performance of the hardware and algorithmic systems. Instead, jointly optimizing the speed constraint and trajectory to optimize (1) can further unlock the potential of the system.
We define a decomposition of the policy $\pi^{\prime}$ which augments $\pi$ in (1) with an auxiliary action $v^{\dagger}$
@@
where $\pi^{\prime}$ is decomposed as two hierarchical policies $\pi^{\prime\prime}$ and $\pi^{\dagger}$, the former of which is conditioned on the auxiliary action while the latter decides it. The action space of the original POMDP is augmented as a result of the introduction of an auxiliary action, i.e., $A^{\prime}=A\times V^{\dagger}$. However, no other part of the POMDP is changed, so we can define a hierarchical policy optimization problem
$\begin{split}\max_{\pi^{\prime\prime},\pi^{\dagger}}\mathbb{E}_{v^{\dagger}_{t}\sim\pi^{\dagger}\left(\cdot|o_{t},g\right),a_{t}\sim\pi^{\prime\prime}(\cdot|v^{\dagger}_{t},o_{t},g)}\left[\sum\nolimits_{t}{\gamma^{t}r\left(s_{t},a_{t},g\right)}\right]=\\ \max_{\pi^{\dagger}}\max_{\pi^{\prime\prime}}\mathbb{E}_{v^{\dagger}_{t}\sim\pi^{\dagger}(\cdot|o_{t},g),a_{t}\sim\pi^{\prime\prime}(\cdot|v^{\dagger}_{t},o_{t},g)}\left[\sum\nolimits_{t}{\gamma^{t}r(s_{t},a_{t},g)}\right],\end{split}$ (5)
equivalent to (1), where, for more concise writing, we omit the state and goal distributions in (1). This problem is equivalent to (1) in the sense that if $\pi$ and $\pi^{\prime}$ impose the same distribution on $a$ when conditioned on any $o\in O$ and $g\in G$, the objective function values of (1) and (5) are consistent.